# The Conspiracy Money Machine: Uncovering

### v2

@@ Abstract
In this paper, we illuminate this ecosystem. First, we propose an approach to detect conspiracy channels. Then, we discover that conspiracy channels can be clustered into four distinct communities comprising over 17,000 channels. Finally, we uncover the "Conspiracy Money Machine," revealing how most conspiracy channels actively seek to profit from their subscribers. We find conspiracy theorists leverage e-commerce platforms to sell questionable products or lucratively promote them through affiliate links. Moreover, we observe that conspiracy channels use donation and crowdfunding platforms to raise funds for their campaigns. We determine that this business involves hundreds of thousands of donors and generates a turnover of almost \$66 million.
@@ 1 Introduction
Conspiracy theories have been an integral part of human history, offering alternative interpretations for complex events [100]. The most common definition, which we use in our work, is that a conspiracy theory is a belief that an event or situation is the result of a secret plan made by powerful people [40]. A notorious example is the Flat Earth theory [36]. Despite centuries of scientific evidence proving the Earth’s roundness, the theory continues to be discussed and promoted by several communities [72, 47]. Throughout history, several conspiracy theories have emerged on a wide range of topics, like the Moon Landing Hoax [46], JFK Assassination [109], Holocaust Denial [66], Elvis Presley’s Faked Death [34], and 9/11 Conspiracy Theories [93].
Nowadays, with the advent of the Internet and social media, conspiracy theories have found new outlets to spread and gain traction [44, 82]. A notable example is the Pizzagate conspiracy theory, which originated and spread on online bulletin boards in 2016 [99]. In 2017, online forums acted as a catalyst for QAnon conspiracy theories [39], which alleged that a global cabal of malevolent elites was involved in heinous activities. The advent of the COVID-19 pandemic has sparked various online conspiracy theories, including claims that the virus is a bio-weapon for population control [56], and that 5G technology is somehow linked to the spread of the virus [8]. Finally, on January 6, 2021, a pro-Trump mob stormed the U.S. Capitol building, disrupting the certification of the 2020 presidential election results [74]. These incidents led the major social media platforms to implement content moderation to curb the dissemination of these theories [18]. In response, conspiracy theorists are flocking to less moderated platforms to freely share their views. Anecdotal evidence from various news sources [20, 103, 4] underscores Telegram, one of the most popular instant messaging applications, as one such platform. This is not surprising, as Telegram offers a permissive content policy and channels—virtual rooms where the admins can broadcast messages to large audiences.
@@
  • Conspiracy Datasets. We release two datasets. The first one is the Conspiracy Resource Dataset, a collection of conspiracy-related web resources gathered through an extensive literature review. The second is the Conspiracy URLs Dataset, a list of 198,818 resolved unique URLs shared by conspiracy theory-related channels. We believe these datasets can enable further studies on identifying and characterizing conspiracy communities on other platforms and determining their activities.
@@
  • Conspiracy Monetization. Finally, we identify the potential strategies that conspiracy theorists can employ to generate revenue from the subscribers of Telegram channels. We find that the most popular approaches involve donations, crowdfunding campaigns, and affiliate programs. We discover 123K URLs linked to donation platforms and more than 24K URLs related to crowdfunding campaigns. Quantifying the amount of money raised with projects sponsored by conspiracy channels, we discovered that they amassed almost \$66M donated by over 906K backers.
@@ 3 Related work
Telegram has recently gained substantial attention, with several prior studies investigating questionable activities on the platform Weerasinghe et al.[107] studied “pods”, organized groups created to artificially boost Instagram popularity. Other studies have explored Telegram’s use in cryptocurrency market manipulations like pump and dump[64, 59] and Ponzi schemes [78]. Additionally, researchers have noted its misuse by terrorist organizations for propaganda and recruitment [27, 110, 108]. Research on conspiracy theories on Telegram is limited, with most studies focusing on other social media platforms. Here is a report on these studies.
@@
4chan/8kun. 4chan [21], an image-based bulletin board, has been linked to conspiracy theories, notably the PizzaGate conspiracy [95]. Papasavva et al.[81] analyzed over 3.5 million messages on 4chan’s /pol board, finding antisemitic conspiracy theories. Similar content is found on 8kun, a platform associated with white supremacism and hate crimes[54, 53]. Aliapoulios et al. [79] discovered that 8kun QAnon threads are significantly larger than those on 4chan.
@@ 4 Methodology
To uncover the structure of channels in Telegram that are related to conspiracy theories, we build a methodology consisting of four steps: First, data collection (including the introduction of a new dataset); second, detection of "conspiracy channels;" third, identification of communities of channels linked to conspiracy theories.
@@ 4.1 Data collection
We leverage two datasets. The first one is the TGDataset [62], the largest collection of public Telegram channels, with over 120,000 channels and 400 million messages. The dataset contains all the messages shared by the collected channels until July 31, 2022. This dataset provides information about the channels (e.g., their title, description, and creation date), all the messages sent with their timestamp and whether a message has been forwarded, and from which channel it originated. Then, we build a novel dataset, the Conspiracy Resources Dataset, that we describe in the following.
4.1.1 Conspiracy Resources Dataset
@@
  • YouTube. We follow the approach of [16] and use two repositories of YouTube channels reported in [65, 33]. These repositories contain a list of 4,007 YouTube channels manually labeled as conspiracy-related. For each channel, we extract the complete list of their videos, resulting in a total of 1,973,439 video IDs.
  • Reddit. We leverage the work of [79, 85, 45, 84] to collect a list of 92 subreddits identified as conspiracy-related.
@@
  • 4chan/8kun. We did not find any resource specifically related to conspiracy theories on 4chan from previous work. Indeed, the infamous /pol discussion board is too general and discusses topics unrelated to conspiracy theories. Instead, we collect a list of 7 boards related to QAnon on 8kun from the work in [79].
  • Websites. From the work of Aliapoulios et al.  [79], we collect 6 websites that are well-known aggregators of Q drops spread by the QAnon conspiracy. Then, we use OpenSources [1] (used also to study QAnon in [54]) to extract 122 website domains linked to conspiracy theories.
Although there are other works mentioning conspiracy theories in other platforms like Parler [9, 17] or Twitter [8], they do not publicly release their data-sets or provide URLs related to conspiracy theories that we can use in our study. Tab 1 reports all the resources we find and the related papers. Moreover, we release in [15] the full dataset of extracted resources.
@@
[65, 16, 33] YouTube 4,007 channels 147,349
[79, 85, 45, 84] Reddit 92 subreddits 17,889
@@ 4.2 Conspiracy channels detection
We devise a methodology that combines the TGDataset and the Conspiracy Resources Dataset to find channels on Telegram related to conspiracy theories. The detection is performed in four steps: First, we extract and pre-process URLs from the TGDataset, then we perform the match with the resources found in the Conspiracy Resources Dataset, we use graph analysis to find clusters of conspiracy-related channels, and finally, we validate our results.
@@ 4.2.1 Data extraction and pre-processing
We parse all the messages (498,320,597) in the 120,979 TGDataset’s channels and use regular expressions to extract all the URLs. In this way, we obtain 205,046,775 URLs, 84,809,578 of which are unique. A first analysis of the URLs reveals that 20.2% (17,140,343 URLs) have been shortened using URL shortener services such as bit.ly (2,368,953 occurrences) or if.tt (1,462,885 occurrences). Since our methodology for detecting conspiracy theories channels revolves around identifying URLs associated with conspiracies, we want to ensure we do not miss any of them because it has been shortened. Thus, between July and September 2023, we resolved the shortened URLs by sending HEAD requests to the shortening service using the Python Requests library [29].
@@ 4.2.2 URLs matching
Then, we extract URLs associated with the resources collected in the Conspiracy Resources Dataset. We start by using the conspiracy channels and videos’ IDs in the Conspiracy Resources Dataset to search YouTube URLs related to conspiracy theories. We detect 2,446 URLs (468 unique) linking to conspiracy channels and 144,903 URLs (58,099 unique) linking to conspiracy videos. The most widely shared channel is Fall Cabal, a channel associated with the "Fall of the Cabal" [7], an antisemitic documentary used to recruit QAnon followers affiliated with Dutch conspiracy theorist Janet Ossebaard [70]. Instead, for Reddit, 8kun, and Voat, we extract all the URLs linking to conspiracy subreddits, boards, and subverses, respectively. We find 17,889 Reddit URLs (17,238 unique), 3,787 8kun URLs (2,542 unique), and 21 Voat URLs (11 unique). Most of the URLs we find are related to the r/conspiracy subreddit, the largest conspiracy theory discussion board on Reddit [85].
Finally, we extract all the URLs having the domain of the flagged websites, finding 299,262 URLs (120,463 unique) from 307 different domains. The website providing the most matches is Zerohedge, a far-right news aggregator known for spreading conspiracy theories, particularly about COVID-19  [75, 76]. The second most popular is InfoWars, well-known for promoting conspiracy theories and fake news [31, 111]. We also detect over 10k URLs linking to the qagg.news website, a popular repository that stores the messages of the QAnon conspiracy theory [49, 79].
In total, we find 468,308 URLs (198,818 unique) posted by 11,713 Telegram channels. We publicly release the extracted URLs in [15] to enable further studies. In the following, we will refer to this dataset as Conspiracy URLs Dataset. Tab. 1 succinctly reports the number of URLs collected inside the Telegram channels divided by resource type.
@@ 4.2.3 Clustering conspiracy channels
The previous analysis shows 11,713 Telegram channels that posted at least one message with a link to a conspiracy-related resource. We study how and if these channels are connected to better understand the phenomenon. To do so, we follow the approach of [63] and build the Telegram forwarding graph for the whole dataset of channels. The forwarding graph is a graph $G=(V,E)$ in which nodes in $G$ are channels and an edge $u\to v$ in $E$ represents the presence in $u$ of a message forwarded from channel $v$. Users of channel $u$ can follow the forwarded message and reach channel $v$. Thus, edges represent the possible flows through channels of users following forwarded messages.
Once we have built the graph, we identify communities—subsets of nodes within the graph that are highly connected with respect to the rest of the graph [86]. In this specific case, a community is a subset of Telegram channels that consistently forward messages among themselves and rarely forward messages from channels of the other communities. To perform community detection, we use the Leiden algorithm [97] since it is widely adopted and has proven to be effective in identifying communities in social graphs (e.g., Twitter [91, 23]). We compute the optimal number of communities with regard to modularity—a metric that measures how much the partitioning is better than a random partition. The metric ranges between 1 to -1, and we obtain a score of 0.78. With this approach, we identify 47 distinct communities of channels. Then, we analyze how the conspiracy-related channels are distributed inside these communities. To illustrate this analysis, we report the scatterplot in Fig. 1. The Figure shows a dot for each community, the number of channels in the community (x-axis), and the percentage of conspiracy channels (y-axis). It is evident that some communities (highlighted in red) stand out due to an unusually high concentration of conspiracy-related channels. In particular, one community has 86% of potential conspiracy channels, while the other 3 have more than 40%. These four communities contain 17,950 channels, almost 15% of all the channels in the TGDataset. Furthermore, a vast fraction (76.6%) of the channels containing at least one link obtained from the Conspiracy Resource Dataset belong to one of these four communities. In the following, we will refer to the channels in the four communities as Conspiracy Channel Dataset and to all the links in these channels as Extended Conspiracy URLs Dataset. Given the relevance of these communities for understanding the diffusion of conspiracy theories on Telegram, we will analyze them in detail.
@@ 4.3 A look into Conspiracy Communities
We adapt this idea to the Telegram graph, defining a channel as authoritative if many good hubs forward its messages, and as a good hub if it forwards messages from many authoritative channels. In our analysis, we are particularly interested in highly authoritative channels. Indeed, according to the definition, these channels are very influential in the community as their messages are widely forwarded. Using the HITS algorithm, we identify and analyze the top five channels with the highest authority scores in each community. Moreover, to better analyze each community, use LangDetect [38] to identify the languages used in channels. In the following, we will analyze separately each community.
English community. This is the largest community, with 9,491 channels and more than 26 million messages sent. More than 89% of these channels communicate using the English language. The most influential channel of this community is Disclose.tv, a website that discusses alternative viewpoints on the news and is notorious for propagating conspiracy theories [89]. Following closely in terms of authority ranking is Tommy Robinson News, a channel allegedly managed by Tommy Robinson, a British activist known to promote conspiracy theories, particularly those related to the threat of Islam for Western societies [35]. RT News is a state-funded international media company headquartered in Russia, known for its alleged bias and for disseminating information that supports the Russian government’s positions [37, 106]. About the Police frequency channel, by searching online, we did not find evidence that this channel is associated with well-known entities or individuals linked to conspiracies. However, looking at its messages, we find that it is a far-right channel that focuses on American news about law enforcement, anti-gun control, and anti-immigration. Finally, Covid Red Pills claims to unveil the truth behind the COVID-19 pandemic.

### v3

@@ Abstract
In this paper, we illuminate this ecosystem. First, we propose an approach to detect conspiracy channels. Then, we discover that conspiracy channels can be clustered into four distinct communities comprising over 17,000 channels. Finally, we uncover the "Conspiracy Money Machine," revealing how most conspiracy channels actively seek to profit from their subscribers. We find conspiracy theorists leverage e-commerce platforms to sell questionable products or lucratively promote them through affiliate links. Moreover, we observe that conspiracy channels use donation and crowdfunding platforms to raise funds for their campaigns. We determine that this business involves hundreds of thousands of donors and generates a turnover of almost \$71 million.
If you cite this paper, please use the USENIX Security reference: Vincenzo Imperati, Massimo La Morgia, Alessandro Mei, Alberto Maria Mongardini, and Francesco Sassi. 2025. "The Conspiracy Money Machine: Uncovering Telegram’s Conspiracy Channels and their Profit Model." In 34th USENIX Security Symposium (USENIX Security 25), pp. 5229-5246. 2025.
https://www.usenix.org/system/files/usenixsecurity25-imperati.pdf
@@ 1 Introduction
Conspiracy theories have been an integral part of human history, offering alternative interpretations for complex events [86]. The most common definition, which we use in our work, is that a conspiracy theory is a belief that an event or situation is the result of a secret plan orchestrated by powerful people or organizations [12, 69]. A notorious example is the Flat Earth theory [32]. Despite centuries of scientific evidence proving the Earth’s roundness, the theory continues to be discussed and promoted by several communities [61, 40]. Throughout history, several conspiracy theories have emerged on a wide range of topics, like the Moon Landing Hoax [39], JFK Assassination [95], Holocaust Denial [86], Elvis Presley’s Faked Death [30], and 9/11 Conspiracy Theories [77].
Nowadays, with the advent of the Internet and social media, conspiracy theories have found new outlets to spread and gain traction [28]. A notable example is the Pizzagate conspiracy theory, which originated and spread on online bulletin boards in 2016 [85]. In 2017, online forums acted as a catalyst for QAnon conspiracy theories [34], which alleged that a global cabal of malevolent elites was involved in heinous activities. The advent of the COVID-19 pandemic has sparked various online conspiracy theories, including claims that the virus is a bio-weapon for population control [47], and that 5G technology is somehow linked to the spread of the virus [6]. Finally, on January 6, 2021, a pro-Trump mob stormed the U.S. Capitol building, disrupting the certification of the 2020 presidential election results [62]. These incidents led the major social media platforms to implement content moderation to curb the dissemination of these theories [15]. In response, conspiracy theorists are flocking to less moderated platforms to freely share their views. Anecdotal evidence from various news sources [17, 90, 4] underscores Telegram, one of the most popular instant messaging applications, as one such platform. This is not surprising, as Telegram offers a permissive content policy and channels—virtual rooms where the admins can broadcast messages to large audiences.
@@
  • Conspiracy Datasets. We release two datasets [80]. The first one is the Conspiracy Resource Dataset, a collection of conspiracy-related web resources gathered through an extensive literature review. The second is the Conspiracy URLs Dataset, a list of 177,567 resolved unique URLs shared by conspiracy theory-related channels. We believe these datasets can enable further studies on identifying and characterizing conspiracy communities on other platforms and determining their activities.
@@
  • Conspiracy Monetization. Finally, we identify the potential strategies that conspiracy theorists can employ to generate revenue from the subscribers of Telegram channels. We find that the most popular approaches involve donations, crowdfunding campaigns, and affiliate programs. We discover 123K URLs linked to donation platforms and more than 24K URLs related to crowdfunding campaigns. Quantifying the amount of money raised with projects sponsored by conspiracy channels, we discovered that they amassed almost \$71M donated by over 906K backers.
@@ 3 Related work
Telegram has recently gained substantial attention, with several prior studies investigating questionable activities on the platform Weerasinghe et al.[94] studied “pods”, organized groups created to artificially boost Instagram popularity. Other studies have explored Telegram’s use in cryptocurrency market manipulations like pump and dump[55] and Ponzi schemes [64]. Additionally, researchers have noted its misuse by terrorist organizations for propaganda and recruitment [96, 75]. Research on conspiracy theories on Telegram is limited, with most studies focusing on other social media platforms. Here is a report on these studies.
@@
4chan/8kun. 4chan [18], an image-based bulletin board, has been linked to conspiracy theories, notably the PizzaGate conspiracy [19]. Papasavva et al.[67] analyzed over 3.5 million messages on 4chan’s /pol board, finding antisemitic conspiracy theories. Similar content is found on 8kun, a platform associated with white supremacism and hate crimes[45]. Papasavva et al. [65] discovered that 8kun QAnon threads are significantly larger than those on 4chan.
@@ 4 Methodology
To uncover the structure of channels in Telegram that are related to conspiracy theories, we build a methodology consisting of three steps: First, data collection (including the introduction of a new dataset); second, detection of "conspiracy channels;" third, identification of communities of channels linked to conspiracy theories.
@@ 4.1 Data collection
We leverage two datasets. The first one is the TGDataset [53], the largest collection of public Telegram channels, with over 120,000 channels and 400 million messages. The dataset contains all the messages shared by the collected channels until July 31, 2022. This dataset provides information about the channels (e.g., their title, description, and creation date), all the messages sent with their timestamp and whether a message has been forwarded, and from which channel it originated. Then, we build a novel dataset, the Conspiracy Resource Dataset, that we describe in the following.
4.1.1 Conspiracy Resource Dataset
@@
  • YouTube. We follow the approach of [12] and use two repositories of YouTube channels reported in [56, 29]. These repositories contain a list of 4,007 YouTube channels manually labeled as conspiracy-related. For each channel, we extract the complete list of their videos, resulting in a total of 1,973,428 video IDs.
  • Reddit. We leverage the work of [65, 70, 38, 69] to collect a list of 91 subreddits identified as conspiracy-related.
@@
  • 4chan/8kun. We did not find any resource specifically related to conspiracy theories on 4chan from previous work. Indeed, the infamous /pol discussion board is too general and discusses as well topics unrelated to conspiracy theories. Instead, we collect a list of 7 boards related to QAnon on 8kun from the work in [65].
  • Websites. From the work of Papasavva et al.  [65], we collect 8 websites that are well-known aggregators of Q drops spread by the QAnon conspiracy. Then, we use OpenSources [1] (also used to study QAnon in [45]) to extract 120 website domains linked to conspiracy theories.
Although there are other works mentioning conspiracy theories in other platforms like Parler [7, 13] or Twitter [6], they do not publicly release their datasets or provide URLs related to conspiracy theories that we can use in our study. Tab 1 reports all the resources we find and the related papers. Moreover, we release in [80] the full dataset of collected resources.
@@
[56, 12, 29] YouTube 4,007 channels 146,915
[65, 70, 38, 69] Reddit 91 subreddits 17,870
@@ 4.2 Conspiracy channels detection
We devise a methodology that combines the TGDataset and the Conspiracy Resource Dataset to find channels on Telegram related to conspiracy theories. The detection is performed in four steps: First, we extract and pre-process URLs from the TGDataset, and we perform the match with the resources found in the Conspiracy Resource Dataset. Then, we use graph analysis to find clusters of conspiracy-related channels, and, finally, we validate our results.
@@ 4.2.1 Data extraction and pre-processing
We parse all the messages (498,320,597) in the 120,979 TGDataset channels and use regular expressions to extract all the URLs. In this way, we obtain 205,046,775 URLs, 84,809,578 of which are unique. A first analysis of the URLs reveals that 20.2% (17,140,343 URLs) have been shortened using URL shortener services such as bit.ly (2,368,953 occurrences) or if.tt (1,462,885 occurrences). Since our methodology for detecting conspiracy theories channels revolves around identifying URLs associated with conspiracies, we want to ensure we do not miss any of them because it has been shortened. Thus, between July and September 2023, we resolved the shortened URLs by sending HEAD requests to the shortening service using the Python Requests library.
@@ 4.2.2 URLs matching
Then, we extract URLs associated with the resources collected in the Conspiracy Resource Dataset. We start by using the conspiracy channels and videos’ IDs in the Conspiracy Resource Dataset to search YouTube URLs related to conspiracy theories. We detect 2,284 URLs (371 unique) linking to conspiracy channels and 144,631 URLs (52,120 unique) linking to conspiracy videos. The most widely shared channel is Fall Cabal, a channel associated with the "Fall of the Cabal" [5], an antisemitic documentary used to recruit QAnon followers affiliated with Dutch conspiracy theorist Janet Ossebaard [59]. Instead, for Reddit, 8kun, and Voat, we extract all the URLs linking to conspiracy subreddits, boards, and subverses, respectively. We find 17,870 Reddit URLs (17,156 unique), 3,768 8kun URLs (2,509 unique), and 21 Voat URLs (11 unique). Most of the URLs we find are related to the r/conspiracy subreddit, the largest conspiracy theory discussion board on Reddit [70].
Finally, we extract all the URLs having the domain of the flagged websites, finding 297,170 URLs (105,877 unique) from 94 different resources. The website providing the most matches is Zerohedge, a far-right news aggregator known for spreading conspiracy theories, particularly about COVID-19  [63]. The second most popular is InfoWars, well-known for promoting conspiracy theories and fake news [97]. We also detect over 10k URLs linking to the qagg.news website, a popular repository that stores the messages of the QAnon conspiracy theory [41, 65].
In total, we find 465,744 URLs (178,044 unique) posted by 11,487 Telegram channels. We publicly release the extracted URLs in [80] to enable further studies. In the following, we will refer to this dataset as Conspiracy URLs Dataset. Tab. 1 succinctly reports the number of URLs collected inside the Telegram channels divided by resource type.
@@ 4.2.3 Clustering conspiracy channels
The previous analysis shows 11,487 Telegram channels that posted at least one message with a link to a conspiracy-related resource. We study how and if these channels are connected to better understand the phenomenon. To do so, we follow the approach of [54] and build the Telegram forwarding graph for the whole dataset of channels. The forwarding graph is a graph $G=(V,E)$ in which nodes in $G$ are channels and an edge $u\to v$ in $E$ represents the presence in $u$ of a message forwarded from channel $v$. Users of channel $u$ can follow the forwarded message and reach channel $v$. Thus, edges represent the possible flows through channels of users following forwarded messages.
Once we have built the graph, we identify communities—subsets of nodes within the graph that are highly connected with respect to the rest of the graph. In this specific case, a community is a subset of Telegram channels that consistently forward messages among themselves and rarely forward messages from channels of the other communities. To perform community detection, we use the Leiden algorithm [83] since it is widely adopted and has proven to be effective in identifying communities in social graphs (e.g., Twitter [76, 20]). We compute the optimal number of communities with regard to modularity—a metric that measures how much the partitioning is better than a random partition. The metric ranges between 1 to -1, and we obtain a score of 0.78. With this approach, we identify 47 distinct communities of channels. Then, we analyze how the conspiracy-related channels are distributed inside these communities. To illustrate this analysis, we report the scatterplot in Fig. 1. The Figure shows a dot for each community, the number of channels in the community (x-axis), and the percentage of conspiracy channels (y-axis). It is evident that some communities (highlighted in red) stand out due to an unusually high concentration of conspiracy-related channels. In particular, one community has 86% of potential conspiracy channels, while the other 3 have more than 40%. These four communities contain 17,950 channels, almost 15% of all the channels in the TGDataset. Furthermore, a vast fraction (78.98%) of the channels containing at least one link obtained from the Conspiracy Resource Dataset belong to one of these four communities. In the following, we will refer to the channels in the four communities as Conspiracy Channel Dataset and to all the links in these channels as Extended Conspiracy URLs Dataset. Given the relevance of these communities for understanding the diffusion of conspiracy theories on Telegram, we will analyze them in detail.
@@ 4.3 A look into Conspiracy Communities
We adapt this idea to the Telegram graph, defining a channel as authoritative if many good hubs forward its messages, and as a good hub if it forwards messages from many authoritative channels. In our analysis, we are particularly interested in highly authoritative channels. Indeed, according to the definition, these channels are very influential in the community as their messages are widely forwarded. Using the HITS algorithm, we identify the top five channels with the highest authority scores in each community, reported in Tab. 2. Moreover, to better analyze each community, use LangDetect [33] to identify the languages used in channels. In the following, we will analyze separately each community.
English community. This is the largest community, with 9,491 channels and more than 26 million messages sent. More than 89% of these channels communicate using the English language. The most influential channel of this community is Disclose.tv, a website that discusses alternative viewpoints on the news and is notorious for propagating conspiracy theories [73]. Following closely in terms of authority ranking is Tommy Robinson News, a channel allegedly managed by Tommy Robinson, a British activist known to promote conspiracy theories, particularly those related to the threat of Islam for Western societies [31]. RT News is a state-funded international media company headquartered in Russia, known for its alleged bias and for disseminating information that supports the Russian government’s positions [93]. About the Police frequency channel, by searching online, we did not find evidence that this channel is associated with well-known entities or individuals linked to conspiracies. However, looking at its messages, we find that it is a far-right channel that focuses on American news about law enforcement, anti-gun control, and anti-immigration. Finally, Covid Red Pills claims to unveil the truth behind the COVID-19 pandemic.